OpenStudy (anonymous):

Where does the equation for simple pendulum motion come from? \[T=2\pi \sqrt{\frac{ l }{ g }}\] where T is period, l is string length, g is acc. due to gravity

OpenStudy (anonymous):

write down the equation of motion for pendulum

OpenStudy (anonymous):



OpenStudy (anonymous):

why is the mg force not straight down?

OpenStudy (anonymous):

sorry...mg is straight down only

OpenStudy (anonymous):

does it help u in finding the time period?

OpenStudy (anonymous):

I want to know how the formula is derived...

OpenStudy (anonymous):

I know write down the equation of motion along the tangent... m( a_t) = mg sin(theta) a_t---tangential acc

OpenStudy (anonymous):

it's fine?

OpenStudy (anonymous):

net force along tangent is mg sin (theta)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

now v_tangential= r ( angular vel)

OpenStudy (anonymous):

differentiate it a_tangential= r(angular acc)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

r is the length of pendulum in this case

OpenStudy (anonymous):

now substitute a_tangential into m( a_t) = mg sin(theta)

OpenStudy (anonymous):

m(r*angular_acc) = m(sin(theta))

OpenStudy (anonymous):

m ( l(angular acc))= mg sin(theta)....l=r= length of pendulum

OpenStudy (anonymous):

now l(angular acc))= g sin(theta)

OpenStudy (anonymous):

where did the l come from?

OpenStudy (anonymous):

l=r=length of pendulum...pendulum is like bob is moving in a circle pivoted at centre

OpenStudy (anonymous):

oh right... l = r. I forgot about that

OpenStudy (anonymous):

now acceleration = g/l sin(theta)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

equation of SHM is a= - w^2 x then time period T= 2 pi/w

OpenStudy (anonymous):

I havent learnt SHM yet. whats w?

OpenStudy (anonymous):

w(omega)---angular frequency acceleration = g/l sin(theta) for small theta acceleration = g/l (theta) sin(theta)= theta

OpenStudy (anonymous):

u know wave... omega= 2 pi (nu)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

so we say SHM is a periodic motion in which acceleration is directed towards a fixed point which is called equilibrium point.

OpenStudy (anonymous):

ok. I know the basics of it

OpenStudy (anonymous):

ok.. so acceleration = g/l (theta)

OpenStudy (anonymous):

for small theta

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

now compare this with a= - w^2 x where x is the displacement from fixed point

OpenStudy (anonymous):

so for small angle x = theta

OpenStudy (anonymous):

so acceleration and displacement are in opposite direction

OpenStudy (anonymous):

yes seems logical

OpenStudy (anonymous):

here theta is x.... for small angle sin(theta)= theta

OpenStudy (anonymous):

so in the case of pendulum angular displacement and angular acceleration are in opposite direction

OpenStudy (anonymous):

write six in the form a power series sin x= x - x^3/3! + x^5/5!-------

OpenStudy (anonymous):

so for small x sin x =x

OpenStudy (anonymous):

neglect higher order terms

OpenStudy (anonymous):

aha

OpenStudy (anonymous):

so we have acceleration =- g/l (theta)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so u have w^2= g/l

OpenStudy (anonymous):

so T= 2 pi/w

OpenStudy (anonymous):

I forgot what w is equal to

OpenStudy (anonymous):

w= angular frequency= 2 pi( frequency)

OpenStudy (anonymous):

so T = 2pi/2pi?

OpenStudy (anonymous):

T= (2 pi)/ (2 pi( frequency))